OpenStudy (anonymous):

the width of a rectangle is 12 units less than the length. the perimeter is 108 units find the length.

OpenStudy (igloo3000):

you do for the length x and width x-12

OpenStudy (igloo3000):

then do x+x+x-12+x-12=108 and add all those up subtract 24 from both sides divide the x's from both sides and you get the answer

OpenStudy (anonymous):

Let the length = x. Then the width should be x-12. To find the perimeter of a rectangle 2(l*w)=Perimeter. So you have a algebraic equation when you sub in your numbers. 2(x*(x-12))=108 Do you understand how I did that?

OpenStudy (igloo3000):

you are actually finding area if you do puddle's way

OpenStudy (anonymous):

Whoops! Sorry. That formula should be 2(l+w)=Perimeter. Thanks igloo :)

OpenStudy (anonymous):

thank you

OpenStudy (igloo3000):

yes another way to say it

OpenStudy (anonymous):

i got 33?

OpenStudy (anonymous):

That's correct :)

OpenStudy (anonymous):

great . thank you (: